OpenStudy (anonymous):

11/12 what kind of number is this? Rational, irrational, integer,etc. I am so confused?

OpenStudy (anonymous):

Aw, what is this, a conspiracy?

jimthompson5910 (jim_thompson5910):

11/12 is a rational number because it is a ratio of two integers.

myininaya (myininaya):

whenever you have or can write you number as integer/integer its rational (and of course the bottom integer is not zero)

myininaya (myininaya):

do you know what an integer is?

hero (hero):

A rational number is a number that is a fraction or one that can be written as a fraction

OpenStudy (anonymous):

rational. all rational numbers have the form A/B. where A and B are real numbers. so all integers are rational, but not all rational numbers are integers

myininaya (myininaya):

here are some examples of an integer: -4,-2,0,1,88888

OpenStudy (anonymous):

are these questions for real? same as mc@xbox

myininaya (myininaya):

here are some of examples of a number that is not an integer: \[\pi, \sqrt{2}. \frac{4}{3}, 0.01\]

myininaya (myininaya):

no naper A/B is rational if A and B are integers

myininaya (myininaya):

B does not equal 0

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Number

OpenStudy (anonymous):

right so for example \[\frac{\pi^2}{6}\]is not rational

myininaya (myininaya):

if you say A/B is rational if A and B are real then you are saying \[\frac{\sqrt{2}}{4}\] is rational

OpenStudy (anonymous):

i think we are all being had

myininaya (myininaya):

which it is not